Suppose that at 100 units of output a monopolist is producing such that marginal revenue is equal to marginal cost. The firm is selling its output at a price of $8 per unit and is incurring average variable costs of $5 per unit and average fixed costs of $4 per unit. On the basis of this information we can conclude that the firm is:
operating at maximum profit by producing the 100 units of output
operating at a loss that could be reduced by shutting down
operating at a profit that could be increased by producing more output
operating at a loss that is less than the loss incurred by shutting down
1
Expert's answer
2014-12-09T09:57:26-0500
If at 100 units of output a monopolist is producing such that MR = MC, and P = $8 per unit, AVC = $5 per unit, AFC = $4 per unit. On the basis of this information, ATC = AVC + AFC = 5 + 4 = $9 per unit, TP = (P - ATC)*Q = (8 - 9)*100 = -$100, but ATC > P > AVC (9 >8 > 5).
So, we can conclude that the firm is: d) operating at a loss that is less than the loss incurred by shutting down.